09:34 - Moved to Run 9 in preparation of moving detectors. No target yet. Beam in the ring at 750 keV/A.
Temperatures OK (attach)
System wide checks OK except 16 fails clock and ADC calib
Rates OK (attach)
Writing to disk OK (attach)
Resetting the sorting script. Moving detectors.
Moved detectors in. See attached. Stripe / shadow quite clear. Some oddities on the top-right DSSD. May be due to ExEy cuts and multiplicity cuts.
Moved on tun run 9 to acquire some stats without motion.
10:13 Temperatures still OK. Rates OK. Leakage currents stable.
We were cutting the beam.
Increased repetition to x5.
Motors fully out - run 10 started. |